Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Please help with veritas P & C example

Expert replies
by sachin_yadav » Sat Dec 11, 2010 11:25 pm
What is the probability that you get a pair when picking the top two cards off of a randomly shuffled deck of cards (52 cards with 13 sets of 4 matching cards) ?


My take was:- 52/52 * 1/51 = 1/51. The question is talking about pairs and asking to pick top two cards, so i believe the pair is of two cards; however this is wrong. The explanation says, this question is asking for the probability to get ANY pair, not one specific pair (say a pair of 2s or 3s). Therefore the probability to get any pair picking the top two cards off the deck is

52/52 * 3/51 = 3/51


Is it a pair of 2 or more ? Please clarify this doubt.
Join the discussion
Source: — Problem Solving |

by kvcpk » Sat Dec 11, 2010 11:55 pm
sachin_yadav wrote:What is the probability that you get a pair when picking the top two cards off of a randomly shuffled deck of cards (52 cards with 13 sets of 4 matching cards) ?


My take was:- 52/52 * 1/51 = 1/51. The question is talking about pairs and asking to pick top two cards, so i believe the pair is of two cards; however this is wrong. The explanation says, this question is asking for the probability to get ANY pair, not one specific pair (say a pair of 2s or 3s). Therefore the probability to get any pair picking the top two cards off the deck is

52/52 * 3/51 = 3/51


Is it a pair of 2 or more ? Please clarify this doubt.
Hi Sachin,

In my explanation, I assume u know cards - symbols.
Let us say we picked K (King-Spade) in the first card.
Now, to get a pair, we can pick up any of these three:
K (King - Clubs)
K (King - Diamonds)
K (King - Hearts)

Hence out of the remaining 51 cards, we have 3 positive scenarios.
Hence probability is 3/51

Hope this helps!!
"Once you start working on something,
don't be afraid of failure and don't abandon it.
People who work sincerely are the happiest."
Chanakya quotes (Indian politician, strategist and writer, 350 BC-275BC)
Join the discussion

by sachin_yadav » Sun Dec 12, 2010 8:12 pm
kvcpk wrote:
Hi Sachin,

In my explanation, I assume u know cards - symbols.
Let us say we picked K (King-Spade) in the first card.
Now, to get a pair, we can pick up any of these three:
K (King - Clubs)
K (King - Diamonds)
K (King - Hearts)

Hence out of the remaining 51 cards, we have 3 positive scenarios.
Hence probability is 3/51

Hope this helps!!
Thanks kvcpk for your reply.

Okay there are 52 cards, i pick one card(king-spade), now to make it a pair i can select 3 cards from the remaining 51 cards . But again there is still one confusion, is it a pair of 2 or 3 ?

What if the question is asking for a pair of 3 ?

So, it will be 52/52 * 2/51. I am not sure.

Waiting for your reply

Please anyone who can tell me the crux of this problem.
Join the discussion

by kvcpk » Sun Dec 12, 2010 9:44 pm
sachin_yadav wrote: Thanks kvcpk for your reply.

Okay there are 52 cards, i pick one card(king-spade), now to make it a pair i can select 3 cards from the remaining 51 cards . But again there is still one confusion, is it a pair of 2 or 3 ?

What if the question is asking for a pair of 3 ?

So, it will be 52/52 * 2/51. I am not sure.

Waiting for your reply

Please anyone who can tell me the crux of this problem.
Couple of things here:
Pair is always 2.
Question specifically mentions "when picking the top two cards off"
Hence we are only looking for a pair of cards which have the same value.

Hope this helps!!
"Once you start working on something,
don't be afraid of failure and don't abandon it.
People who work sincerely are the happiest."
Chanakya quotes (Indian politician, strategist and writer, 350 BC-275BC)
Join the discussion

by sachin_yadav » Mon Dec 13, 2010 12:54 am
kvcpk wrote: Couple of things here:
Pair is always 2.
Question specifically mentions "when picking the top two cards off"
Hence we are only looking for a pair of cards which have the same value.

Hope this helps!!

Thanks once again...
Join the discussion